3.4 Comparing stocks You have the opportunity to invest in either stock A or stock B, and we know the following: Stock A has a standard deviation of 0240, and the correlation between A and the market portfolio is 0.4 Stock B has a standard deviation of 0.3, and the correlation between B and the market portfolio is 0.8 The market portfolio has a variance of 0.04. a) Which stock has the highest beta? b) Which stock has the highest expected return? Motivate your answer using a maximum of 100 words.

hi how do i solve this csn you please explain to someone who have no clue step by step please thank you
Answer #1
a) Beta of a security = SD of the security*COR of the Security with the market/Variance of the market
Beta of A = 0.4*0.4/0.04 = 4
Beta of B = 0.3*0.8/0.04 = 6
STOCK B has the highest beta.
b) Beta is a measure of the systematic risk of a security and the
security with higher beta has higher risk. As systematic risk cannot
be diversified, one has to be compensated for such risk undertaken.
The compensation for risk is the beta times market risk premium
(Market risk premium = Expected market return-Risk free rate).
Hence, expected return of a security under CAPM = risk free rate + beta*expected market risk premium.
Obviously a security having higher beta will have higher expected
return.
Hence, the stock having highest expected return is Stock B.